OpenStudy (anonymous):

1. Show all steps to solve the problem. 5n + 6 = -4 2. Show all steps to solve the problem. 15 = 4a - 5 3. Show all steps to solve the problem. b / -3 + 5 = 7 4. Show all steps to solve the problem. 17 - x / 4 = -10 5. Show all the steps to solve the problem. 3a - 9 = 6. What is the value of 4a - 1? Please help?

OpenStudy (anonymous):

1. 5n+6=-4, subtract 6 from both sides, 5n=-10, divide by 5 to get n by itself, n=-2

OpenStudy (anonymous):

2. 4a-5=15, add 5 to both sides, 4a=20, divide by 4 to get a alone, a=5

OpenStudy (anonymous):

I'm not good at fractions so cant help you out with 3 & 4, sorry. 5. 3a-9=6 add 9 to both sides, 3a=15, divide by 3 to get a alone, a=5

OpenStudy (anonymous):

For 5 substitute the a in 4a-1 with 5. Multiply 4(5)=20-1=19. The value of 4a-1=19

OpenStudy (anonymous):

@ali1029 Thank you so so much! This was very helpful!!

OpenStudy (anonymous):

\[\frac{ b }{ -3 } + 5 =7\] multiply everything by "-3" to get rid of the denominator \[b -15 =-21\] add "15" to each side to get your answer (for question #3) Question #4\[17- \frac{ x }{ 4 }=-10\] multiply everything by "4" to get rid of the denominator:\[68-x=-40\]subtract 68 from each side and solve for "x" but be careful "x' is negative so multiply each side by "-1" after you subtracted 68 from each side.

OpenStudy (anonymous):

you're welcome! :)